OpenStudy (anonymous):

solve using elimination method 5x+2y=-13 7x-3y=17

OpenStudy (saifoo.khan):

x = 0.17 y = -6.06

OpenStudy (anonymous):

3[5x+2y=-13] 2[7x-3y=17] 15x + 6y = - 39 14x - 6y = 34 --------------- 19x = -5 --- --- 19 19 x = -5/19 5x + 2y = -13 5(-5/19) + 2y = -13 -25/19 + 2y = -13 2y = 25/19 - 247/19 2y = -222/19 -- ------- 2 2 y = -222/ 38